OpenStudy (anonymous):

polynomial 4x^2=16 is difference of squares? true or false. and how do you know?

OpenStudy (cwrw238):

rearranging: 4x^2 - 16 = 0 what do you think?

OpenStudy (anonymous):

false?

OpenStudy (cwrw238):

is 16 a square?

OpenStudy (cwrw238):

and 4x^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

yes both of them are 16 = 4^2 and 4x^2 = 2x * 2x and '-' is a difference so its true

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw